• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 16:54:59 on July30,2003

    Entry number 108213

    keyword=sieve in, target collimators out, ice checked

    During controlled access, we inserted the sieve slits and removed the target collimators to get ready for optics run at 2.8 GeV. Also checked ice in the sieve
    box. Again it only was at the lower right corner and not in the acceptance. We
    removed aprt of the ice. Inspected the target ladder, seems all fine.